what is the exact value of sin 22.5 degrees using the sum, difference, half, or double angle formulas and the exact value chart in the solution.
Answer by venugopalramana(3286)

what is the exact value of sin 22.5 degrees using the sum, difference, half, or double angle formulas and the exact value chart in the solution.
LET A = 22.5...LET SIN(A)=X TO BE FOUND OUT..
2A=2*22.5=45
COS(2A)=1-2*{SIN(A)}^2
COS(45)=1-2X^2
1/SQRT(2)=1-2X^2
2X^2=1-1/SQRT(2)=1-SQRT(2)/2=(2-SQRT(2))/2
X^2=(2-SQRT(2))/4
X=SQRT{2-SQRT(2)}/2
=0.3827